How do I get the plot below in R?

95% Prediction Interval Std Dev Fit Area ← 2s Smallest Standard Error
Answer #1

Below is the R code to create this plot:

install.packages("ggplot2")

library(ggplot2)

ggplot(df, aes(x = X, y = Y)) + geom_smooth(method = "lm", se = TRUE) + geom_line(aes(y = Y + 2s), linetype = "dotted") + geom_line(aes(y = Y - 2s), linetype = "dotted")

** If the answer does not match please comment.